In a study to assess the relationship between smoking and CRC in men, Verga Enterprises conducts a study wherein 100 smokers and 100 non-smokers were followed for 10 years (all the men were 60 at the start of the study). Ten men who were smokers developed CRC whereas 5 non-smokers developed disease.

(a) What makes this a Case Control study design?
(b) Complete the 2x2 table.
(c) What is the risk of developing CRC among smokers?
(d) What is the risk of developing CRC among non-smokers?
(e) What is an appropriate measure of association or risk ratio?
(f) What is the risk difference?
(g) What can be concluded about the association between the smoking and CRC?
Answer #1
  1. A case control study is one in which one group which is exposed to a factor is compared with that which isn't ie there is a difference in exposure/outcome. In this case, one group is exposed to cigarette smoke and the second group isn't.
  2. CRC

    no CRC
    Smokers 10 90
    Non-smokers 5

    95

  3. Risk of developing CRC in smokers(Odd's Ratio) = 10*95 / 90*5 = 3.8
  4. Risk in non smokers = 90*5 / 10*95 = 0.47
  5. Odd's Ratio is an appropriate measure of association between exposure and outcome.
  6. The risk difference is 3.8/0.47 = 8.08 (relative risk)
  7. A person who smokes is at 8 times more rush for developing CRC than non smokers.